# Neural radiance fields-based holography [Invited]

### v1

@@ Abstract
This study presents a novel approach for generating holograms based on the neural radiance fields (NeRF) technique. Generating three-dimensional (3D) data is difficult in hologram computation. NeRF is a state-of-the-art technique for 3D light-field reconstruction from 2D images based on volume rendering. The NeRF can rapidly predict new-view images that do not include a training dataset. In this study, we constructed a rendering pipeline directly from a 3D light field generated from 2D images by NeRF for hologram generation using deep neural networks within a reasonable time. The pipeline comprises three main components: the NeRF, a depth predictor, and a hologram generator, all constructed using deep neural networks. The pipeline does not include any physical calculations. The predicted holograms of a 3D scene viewed from any direction were computed using the proposed pipeline. The simulation and experimental results are presented.
@@ 1 Introduction
The calculation of holograms and generation of 3D scenes are also obstacles. Holograms are computed based on a light propagation model and can be categorized into point clouds [12], polygons [13, 14], light fields [15, 16, 17], and deep-learning approaches. The first two represent the 3D scene either as a collection of point cloud data or polygon data. They utilize a point spread function or transfer function representation to calculate light waves from these primitives. Conversely, light-field approaches, including layered holograms, show promise in hologram computation. These approaches automatically render illumination, shading, and hidden surface removal, which are typically cumbersome in point-cloud and polygon approaches. However, this method tends to sacrifice the resolution of the 3D reconstruction. Deep learning has recently emerged as a new hologram-computation technique [18, 19, 20, 21, 22, 23]. These studies have revealed excellent computational performance and image quality; however, all the aforementioned studies require cumbersome and time-consuming 3D scene generation. Several studies proposed the generation of 3D holograms without 3D scene data [24, 25]. These approaches do not require special 3D cameras or computer graphics. These neural networks can predict 3D holograms from 2D images. However, the predicted holograms are only generated from a single viewpoint. In addition, training these neural networks requires a diverse and large 3D dataset for arbitrary 3D reconstruction.
Figure 1: 3D hologram generating pipeline. The first stage performs NeRF to reconstruct a 3D light field from 2D images. The second stage predicts depth maps using MiDaS, and the last stage predicts holograms using Tensor holography.
This study proposes a neural radiance fields (NeRF)-based hologram generation method that can predict holograms directly from new synthesis views without using 3D cameras or 3D graphics processing pipelines. The generation of 3D data is a complicated process in hologram generation. NeRF is a state-of-the-art technique for 3D light-field reconstruction from 2D images [26] that rapidly generates new views that are not included in the training dataset. This study constructs a hologram-rendering pipeline directly converting from a 3D light field generated from 2D images using NeRFs to holograms using deep neural networks, and the pipeline predicts the final holograms within a reasonable time. While conventional hologram calculations involve time-consuming physical calculations, such as diffraction calculations, the pipeline does not include any physical calculations.
@@
2 Prposed pipeline
Figure 1 illustrates the outline of the proposed pipeline. NeRF outputs a new RGB synthesis view by inputting viewing information of the position $(x,y,z)$ and direction angles $(\phi,\theta)$; subsequently, the synthesized image is fed into a deep neural network, MiDaS [27], to predict the corresponding depth map. Concatenating the synthesized RGB image and predicted depth image, it is input to a third deep neural network, “Tensor Holography” [21], to predict a final hologram.
@@ 2 Prposed pipeline
where $\lVert\cdot\rVert_{2}$ denotes $L_{2}$ norm, and $G$ denotes the ground-truth images. In this paper, the loss function is presented in a simplified form. Refere [26] for the actual loss function. This study uses Instant NeRF [27] to accelerate the predictions. In the training, to estimate $(x,y,z,\theta,\phi)$ for each captured image, we used COLMAP software [31]. The captured images and their estimated $(x,y,z,\theta,\phi)$ values were fed into the MLP. Following training, the NeRF can predict novel view images according to the new ${\mathbb{v}}$.
@@
Thus, during the prediction stage, the proposed pipeline requires only ${\mathbb{v}}$. Figure 5 presents the predicted holograms (only the green channel) corresponding to Figs. 3 and 4 using the proposed pipeline. The output holograms are encoded using a double-phase hologram technique [32, 33, 21].
@@ 3.1 Simulations
Figure 6 illustrates the simulated color reconstructions from the holograms predicted in Fig. 5. The angular spectrum method [34] was used for reconstruction. The calculation times for each stage are presented in Fig. 1 are discussed. The most time-consuming part was the optimization of the NeRF. We used 70 captured images, and the optimization time was approximately 1 min to achieve the image quality illustrated in Fig. 3. Following optimization, the NeRF can predict new synthetic images in real-time.
The reconstructed images illustrated in Figs. 6 has low contrast compared to the original images in Fig. 2 due to a domain shift problem in Tensor Holography, which causes the different nature of the data used for training and the data to be predicted. There are two methods to enhance contrast. The first involves transfer learning, wherein Tensor Holography is retrained using our target datasets. Although the retrained neural network predicts the contrast-corrected holograms, it may be necessary to retrain every new target dataset. The second method uses classical contrast enhancement algorithms. Among numerous contrast enhancement algorithms, contrast-limited adaptive histogram equalization (CLAHE) [35] was applied to the holograms predicted from Tensor Holography. In CLAHE, a hologram is divided into nonoverlapping regions, and conventional contrast enhancement is applied to small regions. Figure 7 presents the holograms and histograms of 5 before and after applying CLAHE. The simulated reconstruction of the holograms after applying CLAHE is illustrated in Fig. 8. Thus, the contrast of the reproduced images can be enhanced.
Figure 8: Simulated reconstructions after applying CLAHE to holograms predicted from Tensor Holography.
@@ 3.2 Optical experiments
Optical experiments were conducted as illustrated in Fig. 9. The optical system consisted of an RGB single-mode fiber laser source, lenses, a polarizer, a 4f filter, a spatial light modulator, and a camera. A linear phase pattern was added to the holograms to avoid overlapping of the signal and direct light. The wavelength was set at 520 nm. Figure 10 illustrates the optical reconstruction.
@@
Figure 10: Optical color reconstructions.

### v2

@@ Abstract
This study presents a novel approach for generating holograms based on the neural radiance fields (NeRF) technique. Generating real-world three-dimensional (3D) data is difficult in hologram computation. NeRF is a state-of-the-art technique for 3D light-field reconstruction from 2D images based on volume rendering. The NeRF can rapidly predict new-view images that are not included in a training dataset. In this study, we constructed a rendering pipeline directly from a radiance field generated from 2D images by NeRF for hologram generation using deep neural networks within a reasonable time. The pipeline comprises three main components: the NeRF, a depth predictor, and a hologram generator, all constructed using deep neural networks. The pipeline does not include any physical calculations. The predicted holograms of a 3D scene viewed from any direction were computed using the proposed pipeline. The simulation and experimental results are presented.
@@ 1 Introduction
The calculation of holograms and generation of 3D scenes are also obstacles. Holograms are computed based on a light propagation model and can be categorized into point clouds [12], polygons [13, 14], light fields [15, 16, 17], and deep-learning approaches. The first two represent the 3D scene either as a collection of point cloud data or polygon data. They utilize a point spread function or transfer function representation to calculate light waves from these primitives. Conversely, light-field and layered hologram approaches show promise in hologram computation. These approaches automatically render illumination, shading, and hidden surface removal, which are typically cumbersome in point-cloud and polygon approaches. However, this method tends to sacrifice the resolution of the 3D reconstruction. Deep learning has recently emerged as a new hologram-computation technique [18, 19, 20, 21, 22, 23]. These studies have revealed excellent computational performance and image quality; however, all the aforementioned studies require cumbersome and time-consuming 3D scene generation. Several studies proposed the generation of 3D holograms without 3D scene data [24, 25]. These approaches do not require special 3D cameras or computer graphics. These neural networks can predict 3D holograms from 2D images. However, the predicted holograms are only generated from a single viewpoint. In addition, training these neural networks requires a diverse and large 3D dataset for arbitrary 3D reconstruction.
Figure 1: 3D hologram generating pipeline. The first stage performs NeRF to reconstruct a 3D radiance field from 2D images. The second stage predicts depth maps using MiDaS, and the last stage predicts holograms using Tensor holography.
This study proposes a neural radiance fields (NeRF)-based hologram generation method that can predict holograms directly from new synthesis views without using 3D cameras or 3D graphics processing pipelines. The generation of 3D data is a complicated process in hologram generation. NeRF is a state-of-the-art technique for 3D light-field reconstruction from 2D images [26] that rapidly generates new views that are not included in the training dataset. This study constructs a hologram-rendering pipeline directly converting from a 3D radiance field generated from 2D images using NeRFs to holograms using deep neural networks, and the pipeline predicts the final holograms within a reasonable time. While conventional hologram calculations involve time-consuming physical calculations, such as diffraction calculations, the pipeline does not include any physical calculations.
@@
2 Proposed pipeline
Figure 1 illustrates the outline of the proposed pipeline. NeRF outputs a new RGB synthesis view by inputting viewing information of the position $(x,y,z)$ and direction angles $(\phi,\theta)$; subsequently, the synthesized image is fed into a deep neural network, MiDaS [27], to predict the corresponding depth map. Concatenating the synthesized RGB image and predicted depth image, it is input to a third deep neural network, “Tensor Holography” [21], to predict a final hologram. We have chosen these neural networks that provide state-of-the-art results within a reasonable time.
@@ 2 Proposed pipeline
where $\lVert\cdot\rVert_{2}$ denotes $L_{2}$ norm, and $G$ denotes the ground-truth images. In this paper, the loss function is presented in a simplified form. Refer to [26] for the actual loss function. This study uses Instant NeRF [27] to accelerate the predictions. In the training, to estimate $(x,y,z,\theta,\phi)$ for each captured image, we used COLMAP software [31]. The captured images and their estimated $(x,y,z,\theta,\phi)$ values were fed into the MLP. Following training, the NeRF can predict novel view images according to the new ${\mathbb{v}}$.
@@
Thus, during the prediction stage, the proposed pipeline requires only ${\mathbb{v}}$. Figure 5 presents the predicted holograms (only the green channel) corresponding to Figs. 3 and 4 using the proposed pipeline. The output holograms are encoded using a double-phase hologram technique [32, 33, 21]. Note that in the current pipeline, trained MiDaS and Tensor Holography have been used; only NeRF has been trained.
@@ 3.1 Simulations
Figure 6 illustrates the simulated color reconstructions from the holograms predicted in Fig. 5. The angular spectrum method [34] was used for reconstruction. The calculation times for each stage are presented in Fig. 1 are discussed. The most time-consuming part was the optimization of the NeRF. We used 70 captured images, and the optimization time was approximately 1 min to achieve the image quality illustrated in Fig. 3. The number of images to be input to NeRF is determined empirically from the quality of the output novel view images. Following optimization, the NeRF can predict new synthetic images in real-time. The prediction times of MiDaS and Tensor holography per three color channels were about 5 seconds, respectively. Thus, the proposed pipeline can render holograms viewed from a new view position and an angle of approximately 10 seconds, except for the optimization time of the NeRF. With regard to Tensor Holography, this study used the TensorFlow model; it would be faster with TensorRT.
The reconstructed images illustrated in Figs. 6 has low contrast compared to the original images in Fig. 2 due to a domain shift problem in Tensor Holography, which causes the different nature of the data used for training and the data to be predicted. There are two methods to enhance contrast. The first involves transfer learning, wherein Tensor Holography is retrained using our target datasets. Although the retrained neural network predicts the contrast-corrected holograms, it may be necessary to retrain every new target dataset.
The second method uses classical contrast enhancement algorithms. Denoting a hologram as $h$, a propagation operator as $\mathcal{P}$, the complex amplitude of the reproduced image can be expressed as $u=\mathcal{P}(h)$. Defining a contrast enhancement as $\mathcal{E}$, the complex amplitude of the reproduced image can be expressed as $\mathcal{E}u=\mathcal{P}(\mathcal{E}h)$, and we can observe a contrast-enhanced reproduced image of $|\mathcal{E}u|^{2}$. Among numerous contrast enhancement algorithms, contrast-limited adaptive histogram equalization (CLAHE) [35] was applied to the holograms predicted from Tensor Holography. In CLAHE, a hologram is divided into nonoverlapping regions, and conventional contrast enhancement is applied to small regions. Figure 7 presents the holograms and histograms of Fig. 5 before and after applying CLAHE. The simulated reconstruction of the holograms after applying CLAHE is illustrated in Fig. 8(a). Thus, the contrast of the reproduced images can be enhanced.
Figure 8: Simulated reconstructions after applying CLAHE to holograms predicted from Tensor Holography: (a) corresponding to Fig. 8(a) and (b) corresponding to Fig. 8(b) (Visualizations 1 and 2).
Figure 8(b) shows reproduced images from another hologram using the proposed pipeline. Unlike Fig. 2, this original scene was computer-graphically rendered and new viewpoint holograms were generated from the rendered images.
Note that artifacts in the air are not from depth estimation errors, but from NeRF estimation errors. More images need to be taken for more accurate estimation.
@@ 3.2 Optical experiments
Optical experiments were conducted as illustrated in Fig. 9. The optical system consisted of a single-mode fiber laser source, lenses, a polarizer, a 4f filter, a spatial light modulator, and a camera. A linear phase pattern was added to the holograms to avoid overlapping of the signal and direct light. The wavelength was set at 520 nm. Figure 10 illustrates the optical reconstruction: (a) reconstruction corresponding to Fig. 8(a), and (b) reconstruction corresponding to Fig. 8(b).
@@
Figure 10: Optical reconstructions: (a) reconstruction corresponding to Fig. 8(a) (visualization 3), and (b) reconstruction corresponding to Fig. 8(b) (visualization 4).
@@ 5 Conclusions
[…]
DisclosuresThe authors declare no conflicts of interest.
[…]
Data Availability Statement Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.